Fermi National Laboratory
 

Fermilab Shutdown Report for March 31, 2006 - April 7, 2006

 

LINAC

March 31 – April 7

  • Running to NTF with the H- Ion Source
  • Beam Current 36 ma
  • Brought New Anode Supply On-line for the I- Haefely
  • Characterization the Low Energy RF Systems continues.
  • Low Energy RF PM is Complete.
  • High Energy RF PM continued and is 85% Complete.
  • We continue to Test/Condition Econco Switchtubes in LRF7
  • LRF5 Modulator has failed and troubleshooting continues
  • Replaced Tubes in the H- Haefely

April 7 - 14

  • Continue Characterizing Low Energy RF Systems
  • Finish High Energy RF PM
  • Possibly Install Motor Drive for 10 MeV Collimator
  • Continue Test/Conditioning of Econco Switchtubes LRF7
  • Begin Checking HE RF Phase and Gradient Readback Calibration
  • Install New Digitizers for BPM Readbacks

Linac will begin turning on for 400 MeV beam after the power outages scheduled for April 22 - 24, 2006

Booster

  • MI-8 Alignment nearly complete
  • MI-8 beam line being reconnected
  • Kicker cables pulled
  • Conduits are in – rebuilding MI-8 berm
  • One of two tunnel sump lines completed
  • ORBMP girder may be shipped today
  • Transformer work nearly completed
  • Peek tubing work completed
  • Corrector LCW piping continuing
  • 400 MeV power supply work nearly complete

Schedule Issues

  • MI-8 Dump on schedule – (tasks remaining)
    • Pull and terminate cables
    • After beam tubes connected – begin vacuum work
    • Install kickers after HV testing
    • Septa & Kicker PS connections and testing
  • ORBMP - was scheduled to be here earlier this week
  • 400 MeV line waiting for ORBMP girder install – should begin next week
  • 400 MeV PS work on schedule
  • LCW work on schedule
  • RF work – problems with LCW delaying other jobs

Pbar

On Schedule

Main Injector

  • Continued installation of the new MI8 line multiwires.

  • Continued installation of MI8 collimator.

  • Continued installation of a large aperture quad at 608.
  • Continued installation of large aperture quad at 620.
  • Continued installation of large aperture quad at 101.

  • Began installation of large aperture quad at 402

Recycler

Flying Wire Status

  • Wire Vacuum Certified
  • Tunnel Bake Prep status
    • Disconnection of BPM, Correctors, etc -Done
    • Check bake out circuits -Done
    • Waiting on MI-62 Lambertson Bake to be completed

Damper Kickers

  • Kickers Vacuum Certified (final scan this morning)
  • Tunnel Bake Prep status
    • Bake setup is complete -Done
    • Stands are installed -Done
    • At the stage of cutting into the beam pipe and welding on new flanges

R32 line Vacuum Burst Investigation

  • Replaced Ion pump connector on several ion pumps
  • Prepping the area for a leak checking

Ecool Studies Period

  • Almost doubled the maximum beam current that can be recirculated
    • New ‘world record’
      • 1.8 A, 4.34 MV
      • 7.8 MW of DC beam
  • HV stability being a problem on its own (even at low current), limit to stable operation is not established
  • Max current appears to be limited by the amount of losses to the tubes
    • No indications that the collector is the limitation
  • Shut Down Work
    • Starts next Tuesday April 11th (4-5 weeks)

Ecool Work List

  • New SF6 Skid
  • Repair of the following Vacuum Components
    • IPA03
    • TPS S05
  • Replacement of shaft Bearings
  • Repair Collector Current read back
  • Repair Pulser switch
  • Upgrade Fault protection system
  • Misc. Maintenance

Tevatron

Progress

  • F4 and A3 vacuum work done, walk through complete today
  • A 49 separators installed and under vacuum need to do power test before bake
  • B48 separator installed needs alignment

  • C0 HLS system done
  • D0 HLS system needs final test
  • Shimming on schedule in F sector skipped 11 magnets in E-2 to make way for vacuum work
  • E2 found one bad bellows and one bad seal: finish next week?
  • TEL 2 in place

Vacuum Work

  • F-4 done
  • A-3 done
  • B-4 done
  • E-2 ready for hi pot
  • D-1 leak checking underway D16-1 un rolled
  • A-1 next house to work on

Jobs Not Yet Started

  • Alignment at E-0: xtal Coll, IPM flying wires
  • Re align low beta quads at B0
  • As found A and E sectors with laser tracker
  • A 17 separator
  • TEL 1 installation

SwitchYard

They are ready to install the quadrupole at A0.

NuMI

  • Horn 2 Repair Status
    • Necessary Parts for Repair are Fabricated
    • Practiced Repair Procedure and Finalizing ALARA Plan
    • Expect to Commence with Repair w/o 17-April (2 Days)
  • Target Hall Air Dehumidification – 5 x Desiccant Units
    • Scope of Project Has Management Approval
    • Nearly All Major Requisitions Have Been Submitted
    • Prism Coming Today for Ductwork Estimate
    • Electrical Work Has Started
  • Resealed Target Chase to Allow Fan Restart and Tritium Tests
  • Level Runs in Progress – Currently in Target Hall
  • OTR and Target Profile Monitor Installation has Started
  • Drain Cleaning in Decay Pipe Passageway
  • Target Inspection Not Accomplished

MiniBooNE

  • Horn
    • Ready for safety walkthrough April 13.
    • Interlock enclosure after that.
    • Test week of April 17
  • Alignment
    • Week of May 1?

Collider Detector at Fermilab
On schedule

D-Zero (D0)

  • Layer 0
    • Installed north Layer 0 mount on SMT support structure
    • Prepared for inserting Layer 0 in SMT
      • Configure and align tables/rails and stages
      • Install long tool and guide through SMT while monitoring transverse position
    • Inserted Layer 0 into SMT and mounted on SMT support structure
  • Level 1 Calorimeter Trigger
    • ORC Review
    • Continued connecting BLS cables to patch panels and testing
    • Continued verifying L1Cal Trigger installation
  • Level 1 Central Track Trigger
    • Continued verification of LVDS cabling
  • Detector Maintenance
    • Recovery from power outages for power distribution maintenance
      • VLPC cryostats develop leaks during power outage
      • Readback for L1CTM crate +3.3V supply fails
      • CAL preamp power supply 11 primary fails to recover
    • Continued reassembling individual luminosity monitor counters
    • Continued calorimeter individual channel recovery
    • Upgrades of L1 Muon power supplies in progress
    • Continued muon PDT mods to accommodate latency shift

Mechanical

Work Continues

Cryogenics

Cryo Operations

  • Tevatron at LN2 temperature and LN2 usage is stable with vendor deliveries running only 2-3 truckloads per day due to the Tevatron being 1/4 at room temperature. Helium losses are steady at 15 kscf/day (1/2 of normal during TeV operations).
  • A1, A3, B4, and D1 are at room temperature and isolated with LOTO. E2 and F4 are also at room temperature, have been reconnected by Cryo, and are currently under normal helium flow with anticipation of Division sign-off and cooldown to LN2 temperature soon. F4 house can be cooled down to LN2 temperature without the TEL-1 installed. B4 has been completed by MSD and Cryo will start reconnection procedure. A3 has been completed by MSD, though spool piece conning towers may still be retrofitted (5 out of 8 leak) by Tech Division during this shutdown. Cryo has completed epoxy of helium leaks on Tevatron power leads, though they still need a snoop leak test.
  • Pbar D60 Debuncher vacuum work has been completed by Pbar Dept, thus Cryo personnel will restore helium/nitrogen connections and cool down tanks to LN2 temperature after Division sign-off.
  • ACNET application programs relating to FRIG operations are being tested on Linux consoles.
  • Nitrogen Reliquefier startup is delayed due to concern of “old” electrical cables being used for cross-connect between MSS and KRSS, especially due to high startup currents of 3000 hp TSI compressor. IR Booster compressor was started but had a loading valve failure that is being addressed.
  • MCR monitoring cryo operations.

CHL Maintenance:

  • General facility maintenance is in progress.
  • CHL Helium Compr “A” cooling tower installation has been completed with painting in progress. Some minor controls/electrical/instrumentation work is left.
  • CHL Helium Compr “D” is being overhauled.
  • Helium Coldbox-2 controls upgrade is in progress.

FRIG Maintenance:

  • Ring-wide overhauls of wet expanders in progress with 20/24 completed.
  • Various FRIG bldg vacuum maintenance and instrumentation repairs (e.g. lost charge on VPTs) are in progress, as well as support work on houses at room temperature and other equipment as things fail even during the shutdown period.
  • Various Mycom compressor maintenance: leak hunting and general repairs, F0#1 compressor replacement was completed, F0#3 motor starter replacement in progress.
  • Kautzky change outs have been completed from E4 through F4.
  • The Kautzky vent hose at C4LowBetaQ3 has been lowered to assist survey.

Facility Engineering Services

We finished the work at Casey’s Pond. The west pipe was found half closed due to mussels; the east pipe was a third closed. So far, we’ve collected approximately 9,000 pounds of mussels.

Instrumentation

Work continues.

RF

Work continues.

Controls

Work continues.

ES&H Upgrades

Safety System Upgrades

Keep discussing safety with your crews to keep our focus on getting our jobs done safely.

  • Data Acquisition System 95% Complete
    • 16 of 17 Linux Ethernet Port Replicator (LEPR) chassis installed
  • Booster 50% Complete
    • Booster Dump Move Configuration Changes
    • Completing Documentation to submit for Approval to Operate
  • Main Injector 25% Complete
    • Lower Level Controlled Access Displays Replaced
    • Main Injector Coasting Beam System changes
    • Lam10 Critical Device Interface Installation
  • P-Bar Source 80% Complete
    • Safety System Power Supply Replacement
    • Interlock New BSC700 Beam Stop
  • Tevatron 50% Complete
    • Moving location of both Tevatron Coasting Beam Gate Valves
    • F-Sector Radiation Monitor Card Replacement
  • MT6 Section 1 & 2 70% Complete
    • Enclosure Interlock System Upgrade
  • Safety System Testing 0% Complete
    • All areas need to be re-tested before startup

EE Support

They are working on power supplies.

Alignment

  • Tevatron
    • Completed Unrolls
    • Completed Elevation run (Drops still needed)
    • Completed Deep Rod installation. (Today)
    • Schedule E0 Crystal Collimator
    • Schedule IPM both vert and Horiz
    • Schedule Flying wire cans at E-11.
    • Schedule Flying wires cans at E17
  • Main Injector
    • No Work this week
    • Schedule Next week Deep Rod installation.
  • 8Gev
    • Completed Today The installation of the MI-8 beam dump will require alignment. Several kickers will be installed, one septa magnet, 1 BPM, 2 Toroids, beam pipe, 1 quad will be moved and two multiwires.
    • Scheduled New Texas multiwires will be installed in the MI8 line at 803, 826, 827, 829, and 836. These multiwires will need alignment.
  • P-Bar
    • Completed Elevation runs
    • Completed Set new AP2 beam stop on beam center once it is installed in the Transport enclosure.
  • 400Mev
    • Schedule This week
  • Recycler
    • No work Scheduled
  • MiniBooNE
    • Scheduled May 1
  • NuMI
    • Completed Level run between MI-60 to NuMI SR3 (Tgt Hall) aprox 1500 ft.
  • CDF
    • Completed
  • Linac
    • Completed
  • MIPP
    • Completed
  • Add Ons
    • Tevatron
      • Completed as found Of 5 More additional Rolls
    • Pbar
      • Schedule Topo @ A-50 area next week (Conduit installed)

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/07/2006    email Fermilab